OpenStudy (anonymous):

Someone please help Estimate by first rounding each number to the nearest integer. 1 7/9 - 6 2/11 A. –5 B. –4 C. 4 D. 5

OpenStudy (samanthagreer):

what do you think

OpenStudy (anonymous):

I don't know I need someone to explain how to do this because I didn't get it in my lesson

OpenStudy (samanthagreer):

ok give me one second i will type this all up for you if you ont mind me asking are you home schooled if so what school do you do

OpenStudy (anonymous):

k12

OpenStudy (samanthagreer):

cool me to

OpenStudy (samanthagreer):

B because when you round 1 and 7/9= 2 and 6 and 2/11= 6, and 2-6= negative 4. Answer: B.

OpenStudy (anonymous):

Ok please explain how you rounded those fractions pls

OpenStudy (samanthagreer):

hey here i will go throught it

OpenStudy (samanthagreer):

7/9 is approximately 0.78 2/11 is approximately 0.18 So do we round up or down with 7/9?

OpenStudy (anonymous):

up? I don't know

OpenStudy (samanthagreer):

Yep, So 1 7/9 rounds to 2. What about 2/11?

OpenStudy (anonymous):

down ?

OpenStudy (anonymous):

0 ??

OpenStudy (anonymous):

Wait is it 6?

OpenStudy (samanthagreer):

0.18 If the number after the decimal place is 5 or more, we round up. If it is 4 or less we round down. So do we round up or down?

OpenStudy (samanthagreer):

What's the number after the decimal place(period) 0.18

OpenStudy (anonymous):

18

OpenStudy (samanthagreer):

no 1 and is that 5 or more or is it 4 or less?

OpenStudy (anonymous):

4 or less

OpenStudy (samanthagreer):

Yes, so we round down to 0. So -6 2/11 rounds down to -6. Now subtract 2 - 6

OpenStudy (anonymous):

-4

OpenStudy (samanthagreer):

Yep, that's our answer.

OpenStudy (anonymous):

oh thanks for the help I kind of get it now

OpenStudy (samanthagreer):

great what grade work is this

OpenStudy (samanthagreer):

thanks for the medal

OpenStudy (anonymous):

welcome, and 7th

OpenStudy (samanthagreer):

please romember the close this question

OpenStudy (samanthagreer):

if you need any more help feel free to ask